# Systematic analysis of critical exponents in continuous dynamical phase transitions of weak noise theories

### v1

@@ IV.2 No continuous Fourier frequency components
$\delta I(z)=-\varepsilon\log g_{z}(k_{0},\omega_{0})+\varepsilon\mathcal{C}(k_{\text{max}},\omega_{\text{max}}).$ (30)
@@
$\varepsilon^{\lambda}I_{\text{s}}\left(\delta z/\varepsilon^{\theta}\right)\simeq-\varepsilon\log\left(\delta z/\varepsilon^{\theta}\right),$ (31)
where we have recovered a term $\varepsilon\log\varepsilon^{\theta}$ from $\mathcal{C}$. Setting $\delta z=\varepsilon^{\theta}$ in (31), we find $\lambda=1$. The Ginzburg criterion $\lambda=G_{\mathrm{c}}\theta$ implies
@@ IV.3 Two continuous Fourier frequencies
$\delta I(z)=\varepsilon{\cal C}-\frac{\varepsilon}{\Delta\omega\Delta k}\int_{-k_{\text{max}}}^{k_{\text{max}}}\mathrm{d}k\int_{-\omega_{\text{max}}}^{\omega_{\text{max}}}\mathrm{d}\omega\,\log g_{z\uparrow z_{\mathrm{c}}}(k,\omega)$ (33)
@@ V.1 Large deviations of the spatially averaged height on ring of finite length $\ell$
among the allowed frequencies $k=2\pi m/\ell$ with $m=1,2,3,\dots$ This signals the appearance of a continuous DPT of second order $G_{\text{c}}=2$ [54, 55] at $\bar{H}_{\text{c}}$, the critical exponents of which we will calculate explicitly in the following. Note that, as analyzed in detail in [55], at large $\ell\to\infty$ the second-order DPT at $\bar{H}_{\text{c}}$ is in fact “masked” by another (weakly) first order DPT from the uniform profile to a solitonic solution of (39) which occurs at a smaller $\bar{H}$. Nevertheless, we can still compute the critical behavior of the system in the vicinity of the uniform profile as $\bar{H}\uparrow\bar{H}_{\text{c}}$, even at large $\ell$.
@@
up to terms of order $O\left(\varepsilon^{1/2}\right)$. We let
@@
$I\left(\bar{H}\right)-I_{0}\left(\bar{H}\right)=\delta I\left(\bar{H}\right)\simeq\varepsilon\log\int{\cal D}\left[\delta h\right]\,e^{-\tfrac{1}{2}S_{2}[\delta h]}$ (46)
@@
$\delta I\left(\bar{H}\right)$ $\simeq-\varepsilon\bigg{[}\sum_{\omega\in\pi\mathbb{N}}\log\left(\tfrac{1}{2}g_{\bar{H}}^{(a)}(0,\omega)\right)+\sum_{\begin{subarray}{c}k\in 2\pi\mathbb{N}/\ell\\ \omega\in\pi\mathbb{N}\end{subarray}}\log g_{\bar{H}}^{(a)}(k,\omega)$
@@
$\delta I(\bar{H})\simeq-\varepsilon\bigg{[}\overbrace{\sum_{\omega\in\pi\mathbb{N}}\log\left(\tfrac{\omega^{2}}{4}\right)+\sum_{\begin{subarray}{c}k\in 2\pi\mathbb{N}/\ell\\ \omega\in\pi\mathbb{N}\end{subarray}}\log\left(\tfrac{\omega^{2}}{2}\right)}^{={\cal C}\left(k_{\text{max}},\omega_{\text{max}}\right)}$
@@
for the leading correction in $\varepsilon$ to $I_{0}$ for the uniformly growing profile $\bar{h}$ in the KPZ equation (37) with average height observable (38). It is instructive to compare the result (50) to the one obtained in [54, Eq. (C23)] through a different method, namely solving differential Riccati equations [58, 59, 60] for the SHE. Besides the different choice of units in [54], both results agree when accounting for (i) the cut-off dependent ${\cal C}\left(k_{\text{max}},\omega_{\text{max}}\right)$ here, since the KPZ equation in $1+1$ dimensions, in contrast to the SHE, requires renormalization, and (ii) consequently different $O(1)$ constants that depend on the choice of renormalization and are irrelevant for the critical behavior in the vicinity of $\bar{H}_{\text{c}}$. Regarding the second point, see e.g. [61] for an MFT calculation where, similarly, the cut-off dependent terms are determined through comparison with the exact microscopic result.
@@
The result (50) also confirms the linear stability criterion (40). This criterion gives rise to a complicated structure of the stability region of the uniformly growing interface $\bar{h}(x,t)=\bar{H}t$, which we show in figure 2. The mode $m$ to become unstable first as $\bar{H}$ is increased grows with the domain size $\ell$, with $m\simeq 1.343\ell/2\pi$ and $\bar{H}_{\text{c}}\to 4.603$ as $\ell\to\infty$. At any finite $\ell$, there are either one or two frequencies $k_{0}$ which become unstable at the critical $\bar{H}_{\text{c}}(\ell)$. With regard to the Taylor expansion of $g_{\bar{H}}(k)$ in (23) that determines the CEs, we note that the first derivative of $g_{\bar{H}}(k)$ in (50) at the critical average height and unstable frequencies is given by
@@ V.2 The limit $\ell\to\infty$
$\delta I\left(\bar{H}\right)$ $\simeq-\varepsilon\bigg{[}{\cal C}\left(k_{\text{max}},\omega_{\text{max}}\right)+\tfrac{1}{\Delta k}\int_{0}^{k_{\text{max}}}\mathrm{d}k\;\log g_{\bar{H}}(k)\bigg{]}\,.$ (52)
As argued in section IV.1, even though the exact form of $g_{\bar{H}}(k)$ in (50) is complicated in this example, it suffices to consider its Taylor expansion around $k=k_{0}$ and $\bar{H}=\bar{H}_{\text{c}}(\ell\to\infty)$. The result of this is shown for an example value of $\delta\bar{H}=\bar{H}_{\text{c}}(\ell\to\infty)-\bar{H}=0.1$ in figure 3 and reads $g_{\bar{H}}(k)=A\delta\bar{H}+\tilde{A}(\delta\bar{H})^{2}+B(\delta k)^{2}+\tilde{B}(\delta k)(\delta\bar{H})$ up to second order. Following the calculation of IV.1 shows that the mixed term due to $\tilde{B}\neq 0$ does not change the result, and we still end up with $\lambda=4/3$ and $\theta=2/3$ in the present example since $n=1$.
@@ VI Macroscopic Fluctuation Theory
The MFT is a theoretical framework developed to describe and analyze fluctuations in diffusive nonequilibrium systems. The central idea behind the MFT is to provide a hydrodynamic description of path probabilities, which in turn allow to calculate relevant hydrodynamic observables.
@@
VII Summary
In this work, we have investigated the CEs associated with continuous DPTs in weak noise Langevin systems. Our focus was on the probability of observing atypical events, which we characterized using a large deviation framework. In the context of weak noise Langevin dynamics, the large deviation function is primarily determined by an optimal fluctuation. Close to a DPT, accounting for fluctuations about the optimal fluctuation leads to an anomalous correction to the large deviation function in the form of a singular scaling function (8). By employing the Gaussian fluctuation method introduced here, we classified the CEs $\lambda,\theta$ governing the behavior of this singular scaling function.
We have demonstrated that the CEs $\lambda,\theta$ can be classified into three distinct universality classes, depending on the boundary conditions of the system, when considering a single fluctuating field in $1+1$ dimensions. Extending the analysis to higher dimensions or to systems with multiple fluctuating fields is conceptually straightforward. However, it is important to note that numerical or experimental verification becomes increasingly challenging beyond $1+1$ dimensions.
Although three universality classes of the CEs $\lambda,\theta$ were identified in this work, only two were demonstrated in physical systems. An intriguing future direction would be to explore a weak noise system where the third universality class can be realized. The third universality class requires taking the joint limit of large system size and observation time. This violates the diffusive scaling of the MFT, implying one should look elsewhere for a physical demonstration.
@@ VII Summary
Second, for the $1+1$ dimensional problem studied here, there are two correlation lengths: the spatial correlation length $\xi_{x}$ and the temporal correlation time $\xi_{t}$. We propose that near criticality, $I\sim(\xi_{t}\xi_{x})^{-1}\sim\left(\delta z\right)^{-\nu d_{\perp}}$, where $d_{\perp}$ is the effective number of continuous Fourier modes. Notice that for $d_{\perp}=0$, $\xi_{i}$ with $i\in\{x,t\}$ does not diverge. Otherwise, we recover the scaling relation $2-\alpha=\nu d_{\perp}$. Importantly, the value of $\nu$ may differ from the mean-field exponent $\nu=1/2$, which is valid for the critical dimension $d=d_{\text{c}}=4$ in thermodynamics.
Furthermore, recall that in thermodynamics, the textbook CEs are all connected with empirically measurable quantities. However, in the context of large deviations, it may be more feasible to determine the CEs $\lambda$ and $\theta$ directly through the finite-size corrections to the large deviation function $I$. See [19] for a demonstration of these CEs in a surprisingly small system. In contrast, in systems where the order parameter is non-trivial, it can be challenging to numerically verify the CEs, such as $\beta$, due to the complexities involved in accurately measuring the order parameter’s behavior near criticality, as in [23] where the asymmetry of a one-dimensional KPZ interface with respect to $x=0$ is used as the order parameter for a mirror-symmetry breaking DPT.
@@
Acknowledgements.

### v3

@@ IV.2 No continuous Fourier frequency components
$\delta I(z)=\varepsilon\log g_{z}(k_{0},\omega_{0})+\varepsilon\mathcal{C}(k_{\text{max}},\omega_{\text{max}}).$ (30)
@@
$\varepsilon^{\lambda}I_{\text{s}}\left(\delta z/\varepsilon^{\theta}\right)\simeq\varepsilon\log\left(\delta z/\varepsilon^{\theta}\right),$ (31)
where we have recovered a term $\varepsilon\log\varepsilon^{\theta}$ from $\mathcal{C}$, and a factor of $n$ is omitted as it does not affect the CEs. Setting $\delta z=\varepsilon^{\theta}$ in (31), we find $\lambda=1$. The Ginzburg criterion $\lambda=G_{\mathrm{c}}\theta$ implies
@@ IV.3 Two continuous Fourier frequencies
$\delta I(z)=\varepsilon{\cal C}+\frac{\varepsilon}{\Delta\omega\Delta k}\int_{-k_{\text{max}}}^{k_{\text{max}}}\mathrm{d}k\int_{-\omega_{\text{max}}}^{\omega_{\text{max}}}\mathrm{d}\omega\,\log g_{z\uparrow z_{\mathrm{c}}}(k,\omega)$ (33)
@@ V.1 Large deviations of the spatially averaged height on ring of finite length $\ell$
among the allowed frequencies $k=2\pi m/\ell$ with $m=1,2,3,\dots$ This signals the appearance of a continuous DPT of second order $G_{\text{c}}=2$ [57, 58] at $\bar{H}_{\text{c}}$, the critical exponents of which we will calculate explicitly in the following. Note that, as analyzed in detail in [58], at large $\ell\to\infty$ the second-order DPT at $\bar{H}_{\text{c}}$ is in fact “masked” by another (weakly) first order DPT from the uniform profile to a solitonic solution of (39) which occurs at a smaller $\bar{H}$. Nevertheless, we can still compute the critical behavior of the system in the vicinity of the uniform profile as $\bar{H}\uparrow\bar{H}_{\text{c}}$, even at large $\ell$. We do not address any of the potential critical points in the $\ell$-$\bar{H}$ phase diagram that arise due to the first order transition line here.
@@
up to terms of order $O\left(\varepsilon^{1/2}\right)$. Here, we already omitted the vanishing term $\int_{0}^{1}\mathrm{d}t\int_{0}^{\ell}\mathrm{d}x\;\delta h_{0}\partial_{t}\partial_{xx}\delta h_{0}=-\tfrac{1}{2}\int_{0}^{1}\mathrm{d}t\int_{0}^{\ell}\mathrm{d}x\;\partial_{t}\left(\partial_{x}\delta h_{0}\right)^{2}=0$ in the expansion. We let
@@
$I\left(\bar{H}\right)-I_{0}\left(\bar{H}\right)=\delta I\left(\bar{H}\right)\simeq-\varepsilon\log\int{\cal D}\left[\delta h\right]\,e^{-\tfrac{1}{2}S_{2}[\delta h]}$ (46)
@@
$\delta I\left(\bar{H}\right)$ $\simeq\varepsilon\bigg{[}\sum_{\omega\in\pi\mathbb{N}}\log\left(\tfrac{1}{2}g_{\bar{H}}^{(a)}(0,\omega)\right)+\sum_{\begin{subarray}{c}k\in 2\pi\mathbb{N}/\ell\\ \omega\in\pi\mathbb{N}\end{subarray}}\log g_{\bar{H}}^{(a)}(k,\omega)$
@@
$\delta I(\bar{H})\simeq\varepsilon\bigg{[}\overbrace{\sum_{\omega\in\pi\mathbb{N}}\log\left(\tfrac{\omega^{2}}{4}\right)+\sum_{\begin{subarray}{c}k\in 2\pi\mathbb{N}/\ell\\ \omega\in\pi\mathbb{N}\end{subarray}}\log\left(\tfrac{\omega^{2}}{2}\right)}^{={\cal C}\left(k_{\text{max}},\omega_{\text{max}}\right)}$
@@
for the leading correction in $\varepsilon$ to $I_{0}$ for the uniformly growing profile $\bar{h}$ in the KPZ equation (37) with average height observable (38). It is instructive to compare the result (50) to the one obtained in [57, Eq. (C23)] through a different method, namely solving differential Riccati equations [61, 62, 63] for the SHE. Besides the different choice of units in [57], both results agree when accounting for (i) the cut-off dependent ${\cal C}\left(k_{\text{max}},\omega_{\text{max}}\right)$ here, since the KPZ equation in $1+1$ dimensions, in contrast to the SHE, requires renormalization, and (ii) consequently different $O(1)$ constants that depend on the choice of renormalization and are irrelevant for the critical behavior in the vicinity of $\bar{H}_{\text{c}}$. Regarding the second point, see e.g. [64] for an MFT calculation where, similarly, the cut-off dependent terms are determined through comparison with the exact microscopic result.
@@
The result (50) also confirms the linear stability criterion (40). This criterion gives rise to a complicated structure of the stability region of the uniformly growing interface $\bar{h}(x,t)=\bar{H}t$, which we show in Fig. 2. The mode $m$ to become unstable first as $\bar{H}$ is increased grows with the domain size $\ell$, with $m\simeq 1.343\ell/2\pi$ and $\bar{H}_{\text{c}}\to 4.603$ as $\ell\to\infty$. At any finite $\ell$, there are either one or two frequencies $k_{0}$ which become unstable at the critical $\bar{H}_{\text{c}}(\ell)$. With regard to the Taylor expansion of $g_{\bar{H}}(k)$ in (23) that determines the CEs, we note that the first derivative of $g_{\bar{H}}(k)$ in (50) at the critical average height and unstable frequencies is given by
@@ V.2 The limit $\ell\to\infty$
$\delta I\left(\bar{H}\right)$ $\simeq\varepsilon\bigg{[}{\cal C}\left(k_{\text{max}},\omega_{\text{max}}\right)+\tfrac{1}{\Delta k}\int_{0}^{k_{\text{max}}}\mathrm{d}k\;\log g_{\bar{H}}(k)\bigg{]}\,.$ (52)
As argued in section IV.1, even though the exact form of $g_{\bar{H}}(k)$ in (50) is complicated in this example, it suffices to consider its Taylor expansion around $k=k_{0}$ and $\bar{H}=\bar{H}_{\text{c}}(\ell\to\infty)$. The result of this is shown for an example value of $\delta\bar{H}=\bar{H}_{\text{c}}(\ell\to\infty)-\bar{H}=0.1$ in Fig. 3 and reads $g_{\bar{H}}(k)=A\delta\bar{H}+\tilde{A}(\delta\bar{H})^{2}+B(\delta k)^{2}+\tilde{B}(\delta k)(\delta\bar{H})$ up to second order. Following the calculation of IV.1 shows that the mixed term due to $\tilde{B}\neq 0$ does not change the result, and we still end up with $\lambda=4/3$ and $\theta=2/3$ in the present example since $n=1$.
To further cement the predictions of the Gaussian fluctuation method, we now turn to a different set of hydrodynamic theories – the macroscopic fluctuation theory. We will recall the results of [19], and recast them in the language of the Gaussian fluctuation method.
@@ VI Macroscopic Fluctuation Theory
The MFT is a theoretical framework developed to describe and analyze fluctuations in diffusive and interacting nonequilibrium systems. The central idea behind the MFT is to provide a hydrodynamic description of path probabilities, which in turn allow to calculate relevant hydrodynamic observables.
@@
VII Summary and discussion
In this work, we have investigated the CEs associated with continuous DPTs in weak noise Langevin systems. Our focus was on the probability of observing atypical events, which we characterized using a large deviation framework. In the context of weak noise Langevin dynamics, the large deviation function is primarily determined by an optimal fluctuation. Close to a DPT, accounting for fluctuations about the optimal fluctuation leads to an anomalous correction to the large deviation function in the form of a singular scaling function (8). By employing the Gaussian fluctuation method introduced here, we classified the CEs $\lambda,\theta$ governing the behavior of this singular scaling function. Previously, identifying the CEs in weak noise theories was possible mainly due to the identification of an order parameter, and developing a Landau theory, suggesting that only a limited set of problems could be handled. The Gaussian fluctuation method introduced here offers a broader approach to identifying the CEs $\lambda,\theta$.
We have demonstrated that the CEs $\lambda,\theta$ can be classified into three distinct universality classes, depending on the boundary conditions of the system, when considering a single fluctuating field in $1+1$ dimensions. Extending the analysis to higher dimensions or to systems with multiple fluctuating fields is conceptually straightforward. For higher spatial dimensions, the Fourier modes $k$ become vectors. The task reduces once again to count the number of required continuous Fourier modes which leads to more possibilities. One would need to perform higher dimension integrals, like (33), to infer the CEs. For the case of additional fluctuating fields, one would express them using the Fourier decomposition as well, and obtain matrix-valued coefficient functions $g_{z}(k,\omega)$. Typically, one expects that at the critical point, it is only one of the eigenvalues of $g_{z}(k_{0},\omega_{0})$ that becomes zero, thus recovering our previous treatment.
However, it is important to note that numerical or experimental verification becomes increasingly challenging beyond $1+1$ dimensions.
In this work, we have demonstrated the Gaussian fluctuation method of classifying the CEs for two cases: DPTs in the short-time KPZ equation and diffusive systems within the framework of the MFT. One may ask whether the Gaussian fluctuation method could still be valid for theories with long-range interactions where the action may become non-local [28], or to include memory effects where additional fluctuating fields may be present [80]. For the case of additional fluctuating fields was already discussed, and the theory remains intact. For non-local actions, the theory still applies as well. However, providing an tractable example where one can still fully diagonalize the Gaussian fluctuations, i.e. find $g_{z}(k,\omega)$ explicitly, will be particularly challenging due to the non-locality of the action. Hence, demonstrating the Gaussian fluctuation method for such a system remains a future challenge. Finally, we do not expect the Gaussian fluctuation theory to hold for BKT phase transitions.
Although three universality classes of the CEs $\lambda,\theta$ were identified in this work, only two were demonstrated in physical systems. An intriguing future direction would be to explore a weak noise system where the third universality class can be realized. The third universality class requires taking the joint limit of large system size and observation time. This violates the diffusive scaling of the MFT, implying one should look elsewhere for a physical demonstration. Additionally, all examples considered in this work involve continuous DPTs of order $G_{\text{c}}=2$, with $n=1$ for the unstable frequency coefficient in (23). Finding and studying physical systems with other values of these parameters, perhaps even with vanishing denominators in (29) or (36), would be interesting.
@@ VII Summary and discussion
Second, for the $1+1$ dimensional problem studied here, there are two correlation lengths: the spatial correlation length $\xi_{x}$ and the temporal correlation time $\xi_{t}$ associated with the dynamical exponent. We propose that near criticality, $I\sim(\xi_{t}\xi_{x})^{-1}\sim\left(\delta z\right)^{-\nu d_{\perp}}$, where $d_{\perp}$ is the effective number of continuous Fourier modes. Notice that for $d_{\perp}=0$, $\xi_{i}$ with $i\in\{x,t\}$ does not diverge. Otherwise, we recover the scaling relation $2-\alpha=\nu d_{\perp}$. Importantly, the value of $\nu$ may differ from the mean-field exponent $\nu=1/2$, which is valid for the critical dimension $d=d_{\text{c}}=4$ in thermodynamics. We note that depending on the continuous spatial/temporal degrees of freedom, the correlation scales $\xi_{x,t}$ either diverge with exponent $\delta z^{-\nu}$ or remain finite.
Furthermore, recall that in thermodynamics, the textbook CEs are all connected with empirically measurable quantities. However, in the context of large deviations, it may be more feasible to determine the CEs $\lambda$ and $\theta$ directly through the finite-size corrections to the large deviation function $I$. See [19] for a demonstration of these CEs in a surprisingly small system. In contrast, in systems where the order parameter is non-trivial, it can be challenging to numerically verify the CEs, such as $\beta$, due to the complexities involved in accurately measuring the order parameter’s behavior near criticality, as in [23] where the asymmetry of a one-dimensional KPZ interface with respect to $x=0$ is used as the order parameter for a mirror-symmetry breaking DPT. We also comment that while an experimental study of large deviations and dynamical phase transitions seem challenging, several works have recently made advances into observation of large deviations in general [81], and DPTs in particular [82].
@@
As a final remark, we have assumed throughout this work that the hydrodynamic theory encapsulates the range of possible phenomena, and furthermore that microscopic agents cannot affect its behavior, leading to independence in choosing microscopic cutoffs. While this is generally true, one can point out cases where these assumptions break down, e.g. condensation of particles on a microscopic length scale [90], shock waves in gases due to a moving piston [91], and effects of a local battery drive [92, 93]. It would be interesting to develop a rigorous criterion, revealing when hydrodynamic theories are self-consistent, such that microscopic cutoffs cannot alter the macroscopic behavior.
Acknowledgments.